OpenStudy (anonymous):

Hey, for HW #5 question #4: for part (a) are we supposed to use the end-effector configuration that Dr. Vela just went over in class where you have ge(alpha) where alpha is a vector and then we define R(alpha) and d(alpha). In the notes these end up being really big matrices. also for part (b) would we just have ge(alpha)=g1(alpha1)g2(alpha2)g3(alpha3)g4(alpha4) ?

OpenStudy (anonymous):

I have another question pertaining to the same problem. If class when we went over this Dr. Vela drew the zero-configuration picture and that is what he used to determine his g1(aplha), g2(alpha2), etc. Are we to do the same thing for this problem? For instance for g2(alpha2) I am getting R(alpha2) = \[\left[\begin{matrix}\cos(\alpha2) & -\sin(\alpha2) \\ \sin(\alpha2) & \cos(\alpha2)\end{matrix}\right]\] and also d(alpha2) =\[\left[ 0 0 l1 \right]^{T}\] I don't know if that is right or not, but I'm not sure what else to do.

OpenStudy (anonymous):

I think you are looking at the more complex example in 3d space, problem 4 is only in 2d space. the d(alpha2) should be pretty simple = [l1, 0]T

OpenStudy (anonymous):

Let me try to answer one by one...For part A you should be doing some geometry. It's exactly the same thing as what was worked out in the first week. It should have some sines, cosines, l1, l2, l3, etc. I think what you're thinking of is the Jacobean matrix which does not apply here. For part B what you have is correct but you should be more specific. you should write out each g matrix, including R and d components. As a hint, the first g has a rotation by alpha1 and no displacement. Your R for alpha2 looks right but the d, as the previous answer pointed out, should be \[\left(\begin{matrix}l2 \\ 0\end{matrix}\right)\] Does that clear it up any?

OpenStudy (anonymous):

Thanks for your help! That does clear things up. I still have a question about the d for alpha2. Why is it \[\left(\begin{matrix}l2 \\ 0\end{matrix}\right)\] and not \[\left(\begin{matrix}l1 \\ 0\end{matrix}\right)\]

OpenStudy (anonymous):

My apologies. Your answer is correct.

OpenStudy (anonymous):

Oh ok thanks!

OpenStudy (anonymous):

Why is d not [ l1cos(alpha1), l1sin(alpha1)] ?

OpenStudy (anonymous):

Because remember that in the first transformation, you are rotating by alpha1. The next transformation should be in relation to the new coordinate axes, so in this case only along the x axis. \[g^0_{\alpha2} = g^0_{\alpha 1}*g^{\alpha 1}_{\alpha 2}\]So the second g in your kinematic chain should be the transformation from alpha1 to alpha2, not all the way from the beginning.

OpenStudy (anonymous):

OK that makes sense. Would the rotation at alpha4 would have a d = {0 0} ?

OpenStudy (anonymous):

You don't need to take into account the distance to the end effector frame but g(alpha4) does have a displacement.

OpenStudy (anonymous):

In the notes, for the 3D configuration, my end effector chain ends in g5 ge(alpha) = g1(alpha1)*g2(alpha2)*g3(alpha3)*g4(alpha4)*g5 g5 was basically an identity matrix with an l3 included to adjust for the length of the end effector arm. Since this, 2D, problem has no end effector arm length, I would have the g5 (equivalent) to be an indentity matrix. Am I thinking ablout this correctly?

OpenStudy (anonymous):

I think in the example, g5 is just a length from the final joint to the end effector. In this case, the end effector can rotate and is a joint in itself. You can treat eat as any other joint. The g for alpha4 is analogous to the g for alpha5.

OpenStudy (anonymous):

alpha3 is what I meant at the end there

OpenStudy (anonymous):

So, if my configuration ends in a final rotation with no length to the end effector then my end effector chain will end in a g(alpha) ? ge(alpha) = g1(alpha1)*g2(alpha2)*g3(alpha3)

OpenStudy (anonymous):

no length from the joint to the end effector

OpenStudy (anonymous):

*g4(alpha4). The last joint is the end effector essentially. How did you define the g for alpha3?

OpenStudy (anonymous):

I wrote the last ge(alpha) equation as an example. For this problem, my g3(alpha3) is d = {l2 0} R = [ cos(alpha3) -sin(alpha3) ] [ sin(alpha3) cos(alpha3)]

OpenStudy (anonymous):

Ok that's right. The last joint is exactly analogous to that right? So you still need to make g4(alpha4) , but you don't have to do a g5 to get to the end effector you can just end at g4.

OpenStudy (anonymous):

For 4a, I believe we did something like this in a prior homework. In that homework there was no rotation joint at the end effector therefore no alpha4. Does not alpha4 need to be taken into account for 4a?

OpenStudy (anonymous):

Yes it does because there is a rotation at the end effector joint. That's why you need g4(alpha4)

OpenStudy (anonymous):

In working through problem 4, do you need to do 4b before 4a?

OpenStudy (anonymous):

Oh I see what you're talking about...alpha4 doesn't affect the location of the end effector but it affects the orientation.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes so in 4a for your expression of the location of the end effector, alpha4 does not come into play but in your expression for the angle it does.

OpenStudy (anonymous):

I would have a d and an R expression for 4a that is the result of 4b?

OpenStudy (anonymous):

No you would have an expression for x_e, y_e, and theta_e. It's the same thing that was done the first week of class using geometry.

OpenStudy (anonymous):

4a should not use a kinematic chain of Lie groups. It's purely geometry.

OpenStudy (anonymous):

I feel comfortable using geometry until I get to alpha4 where no length is involved.

OpenStudy (anonymous):

Let me work on it a moment

OpenStudy (anonymous):

Alpha4 does not affect x_e or y_e. What is your expression for theta? Hint: don't overthink it.

OpenStudy (anonymous):

I do not have a theta.

OpenStudy (anonymous):

Just intuitively what should be the angle of the end effector?

OpenStudy (anonymous):

I woud say the angle of the end effector should be the sum alpha1 + alpha2 + alpha3 but the end effector would not close on the object as desired.

OpenStudy (anonymous):

That's close by why would alpha4 not get added on to that?

OpenStudy (anonymous):

xe = l1cos(alpha1) + l2cos(alpha1 + aplha2) + l3cos(alpha1 + aplha2 + alpha3) where would I add alpha4?

OpenStudy (anonymous):

Again, x_e and y_e are not affected at all by alpha4. It's only a rotation and therefore only affects the angle of the end effector, not the location.

OpenStudy (anonymous):

I would not have an R because I am only using geometry not kinematics. x_e and y_e are location values that are arrived at through geometry. I need to use alpha4 for rotation only. I understand these statements but I do not see how to put the pieces together.

OpenStudy (anonymous):

You're really overthinking it....Your x_e is right. You may not have an "R", but you still have an angle of the end effector, which in this case equals alpha1+alpha2+alpha3+alpha4.

OpenStudy (anonymous):

I do understand that the answer would involve theta = alpha1 + alpha2 + alpha3 + alpha4 If I multiply x_e by cos(theta), I know that is not correct If I add cos(theta) to x_e, I know that makes little sense I can not add alpha4 to the last cos factor xe = l1cos(alpha1) + l2cos(alpha1 + aplha2) + l3cos(alpha1 + aplha2 + alpha3 + alpha4) that would out me at a different location with xe = l1cos(alpha1) + l2cos(alpha1 + aplha2) + l3cos(alpha1 + aplha2 + alpha3) I have reached the end effector, and now I need to turn so my next move will be something in addition to my last moves. something that turns the end effector.

OpenStudy (anonymous):

Ok let me just summarize from the beginning: You are looking for the following representation\[\left(\begin{matrix}x_e \\ y_e \\ \theta e\end{matrix}\right)\]x_e and y_e are not affected by alpha4. You wrote the correct equation for x_e above. I'm sure you know how to write the one for y_e also. The only thing alpha4 affects is theta_e. It affects in a simple way - it is added to the other alphas to figure out the final angle of the end effector such that\[\theta_e = \alpha_1+\alpha_2+\alpha_3+\alpha_4\] That's all there is to it.

OpenStudy (anonymous):

I thought I was looking for \[\left(\begin{matrix}x _{e} \\ y _{e}\end{matrix}\right)\] not \[\left(\begin{matrix}x _{e} \\ y _{e} \\ \theta \end{matrix}\right)\]

OpenStudy (anonymous):

That is not correct. As with any frame in SE2, it needs a location and a rotation to completely describe it.

OpenStudy (anonymous):

Ok thank you for all your help

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

in the comment section of the Adjoint.m code, the adjoint id defined as Ad_g1 (g2) = g1 * g2 * inverse g1 should it not be Ad_g1 (g2) = inverse g1 * g2 * g1

OpenStudy (anonymous):

Thanks pvpman